### v2

Abstract
In this paper, we consider a sensor placement problem where sensors can move within a network over time. Sensor placement problem aims to select $K$ sensor positions from $N$ candidates where $K<N$. Most existing methods assume that sensor positions are static, i.e., they do not move, however, many mobile sensors like drones, robots, and vehicles can change their positions over time. Moreover, underlying measurement conditions could also be changed, which are difficult to cover with statically placed sensors. We tackle the problem by allowing the sensors to change their positions in their neighbors on the network. We sequentially learn the dictionary from a pool of observed signals on the network based on graph sampling theory. Using the learned dictionary, we dynamically determine the sensor positions such that the non-observed signals on the network can be best recovered from the observations. In experiments, we validate the effectiveness of the proposed method via the mean squared error (MSE) of the reconstructed signals. The proposed dynamic sensor placement outperforms the existing static ones for both synthetic and real data.
{IEEEkeywords}
Dynamic sensor placement on graphs, dictionary, sparse coding, graph sampling theory
1 INTRODUCTION
[…]
Sensor networks are used in various engineering applications such as traffic, infrastructure, and facility monitoring systems [1, 2, 3, 4]. They aim to efficiently collect information from distributed sensor measurements to enable real-time monitoring and analysis. Sensing targets include many environmental and/or behavioral data like temperature and acceleration [5, 6].
Sensor placement problem is one of the main research topics in sensor networks [7, 8, 9]. Its purpose is to find $K$ sensor positions among $N$ candidates $(K<N)$ where sensors can be placed optimally so that the reconstruction error is minimized in some sense. It has been extensively studied in machine learning and signal processing [10, 11, 8, 12].
In many cases, sensors can be assumed to form a network, including wireless sensor networks, the Internet, and power grids. Therefore, sensor placement on a network, i.e., graph, has been widely studied [13]. In the graph-based sensor placement problem, $N$ nodes of the graph represent possible sensor positions, the edges correspond to relationships among the sensors, and the signal value on a node is a measurement.
Many existing works of sensor placement problem assume the static sensor placement, i.e., sensors do not or cannot move [14, 15, 16]. However, the measurement conditions could be changed over time. In this case, the selected sensors in the static positions do not reflect the current signal statistics. This may lead to the limitation of reconstruction qualities.
To overcome the limitation of the above-mentioned static sensor placement, we consider a dynamic sensor placement problem where sensors can flexibly change their positions over time. Typically, we consider a dynamic placement: sensors that can move dynamically like drones and robots.
To realize the dynamic sensor placement, we sequentially learn the dictionary for signal reconstruction from a pool of observed signals on a network (graph signals) by utilizing sparse coding. We then dynamically determine the sensor positions at every time instance such that the non-observed graph signal can be best recovered from the sampled graph signal.
In experiments, we demonstrate that the proposed method outperforms existing sensor placement methods both in synthetic and real datasets.
Notation: Symbols used in this paper are summarized in Table 1. We consider a weighted undirected graph $\mathcal{G}=(\mathcal{V,E})$, where $\mathcal{V}$ and $\mathcal{E}$ represent sets of nodes and edges, respectively. The number of nodes is $N=|\mathcal{V}|$ unless otherwise specified. The adjacency matrix of $\mathcal{G}$ is denoted by $\mathbf{W}$ where its $(m,n)$-element $\mathbf{W}_{mn}\geq 0$ is the edge weight between the $m$th and $n$th nodes; $\mathbf{W}_{mn}=0$ for unconnected nodes. The degree matrix $\mathbf{D}$ is defined as $\mathbf{D}=\text{diag}\,(d_{0},d_{1},\ldots,d_{N-1})$, where $d_{m}=\sum_{n}\mathbf{W}_{mn}$ is the $m$th diagonal element. We use graph Laplacian $\mathbf{L}\coloneqq\mathbf{D}-\mathbf{W}$ as a graph variation operator. A graph signal $\mathbf{x}\in\mathbb{R}^{N}$ is defined as a function $x:\mathcal{V}\rightarrow\mathbb{R}$. Simply speaking, $x[n]$ is located on the node $n\in\mathcal{V}$ of $\mathcal{G}$.
The graph Fourier transform (GFT) of $\mathbf{x}$ is defined as
$\hat{x}(\lambda_{i})=\braket{\mathbf{u}_{i},\mathbf{x}}=\sum_{n=0}^{N-1}u_{i}[n]x[n],$ (1)
where $\mathbf{u}_{i}$ is the $i$th column of an orthonormal matrix $\mathbf{U}$. It is obtained by the eigendecomposition of the graph Laplacian $\mathbf{L}=\mathbf{U\Lambda U}^{\top}$ with the eigenvalue matrix $,\mathbf{\Lambda}=\mathrm{diag}\,(\lambda_{0},\lambda_{1},$$\ldots,\lambda_{N-1})$. We refer to $\lambda_{i}$ as a graph frequency.
2 GRAPH SIGNAL SAMPLING
In this section, we introduce graph signal sampling. We first review sampling theory on graphs and then describe a sampling set selection on graphs in [17].
2.1 SAMPLING THEORY ON GRAPHS
Here, we assume graph signals (i.e., sensor measurements) inherit the underlying graph information like smooth signals on a network dominant at low graph frequency [18]. Suppose that a graph signal is characterized by the following model [19]:
$\mathbf{x}\coloneqq\mathbf{A}\mathbf{d},$ (2)
[…]
Table 1: Symbols used in this paper.
$N$ Number of nodes
$K$ Number of sensors
$T$ Time duration (date length)
$D$ Number of time slots
$t$ Index of the time instance
[…]
$\|\cdot\|_{1}$ $\ell_{1}$ norm
$\|\cdot\|_{F}$ Frobenius norm
$\mathbf{Y}_{:i}$ $i$th column of $\mathbf{Y}$
$\mathbf{Y}_{mn}$ $(m,n)$-th element of $\mathbf{Y}$
$\mathbf{Y}_{\mathcal{XY}}$ Submatrix of $\mathbf{Y}$ indexed by $\mathcal{X}$ and $\mathcal{Y}$
$\mathbf{Y}_{\mathcal{X}}$ $\mathbf{Y}_{\mathcal{XX}}$
$\mathcal{X}^{c}$ Complement set of $\mathcal{X}$
$\mathcal{R}(\mathbf{Y})$ Range space of $\mathbf{Y}$
The bandlimited (BL) model is one of the well-studied generators [19, 20, 21, 22, 23, 24], defined as
$\mathbf{x}=\sum_{i=0}^{M-1}d[i]\mathbf{u}_{i}=\mathbf{U}_{\mathcal{VM}}\mathbf{d},$ (3)
where $\mathbf{U}_{\mathcal{VM}}\in\mathbb{R}^{N\times M}$ is the submatrix of $\mathbf{U}$ whose columns are specified by =$\mathcal{M}=\{1,\ldots,M\}$. In this case, the generator matrix is $\mathbf{A}=\mathbf{U}_{\mathcal{VM}}$ where $M$ is the bandwidth.
Piecewise constant (PC) model [25] is another possible signal model and is defined as follows.
$\mathbf{x}=\sum_{i=0}^{M-1}d[i]\mathbf{1}_{\mathcal{T}_{i}}=[\mathbf{1}_{\mathcal{T}_{1}},\dots,\mathbf{1}_{\mathcal{T}_{M}}]\mathbf{d},$ (4)
where $\mathcal{T}_{i}\,(i=1,\ldots,M)$ is the nonoverlapping subset of nodes for the $i$th cluster: $[\mathbf{1}_{\mathcal{T}_{i}}]_{n}=1$ if the node $n$ is in $\mathcal{T}_{i}$ and 00 otherwise [25]. In this case, the generator matrix is $\mathbf{A}=[\mathbf{1}_{\mathcal{T}_{1}},\dots,\mathbf{1}_{\mathcal{T}_{M}}]$ where $M$ is the number of clusters.
If we do not know the exact $\mathbf{A}$ but have its prior knowledge, it may be learned from the observations. Let $\mathbf{X}\in\mathbb{R}^{N\times M}$ be a collection of $M$ graph signals. Typically, $\mathbf{A}$ has been determined by $\mathbf{A}=\bm{\mathcal{U}}$, where $\bm{\mathcal{U}}$ is obtained by the SVD of the collection of observed signals $\mathbf{X}\coloneqq\bm{\mathcal{U}}\mathbf{\Sigma}\bm{\mathcal{V}}^{\top}$ [26], and $\bm{\mathcal{U}}$ and $\bm{\mathcal{V}}$ are unitary matrices.
[…]
$\tilde{\mathbf{x}}=\mathbf{A}(\mathbf{S}^{\top}\mathbf{A})^{\dagger}\mathbf{c},$ (5)
where $\cdot^{\dagger}$ is the Moore-Penrose pseudoinverse.
If $\mathcal{A}$ and $\mathcal{S}$ together span $\mathbb{R}^{N}$ and only intersect at the origin, perfect recovery, i.e., $\tilde{\mathbf{x}}=\mathbf{x}$, is guaranteed. This is referred to as the direct sum (DS) condition [19]. Note that the DS condition implies whether $\mathbf{S}^{\top}\mathbf{A}\mathbf{A}^{\top}\mathbf{S}$ is invertible, since $(\mathbf{S}^{\top}\mathbf{A})^{\dagger}=\mathbf{A}^{\top}\mathbf{S}(\mathbf{S}^{\top}\mathbf{A}\mathbf{A}^{\top}\mathbf{S})^{-1}$ holds if and only if the DS condition is satisfied.
From the perspective of the sensor placement, $\mathbf{S}^{\top}$ can be considered as a node-wise sampling operator. In the following, we introduce a design method of $\mathbf{S}^{\top}$ on $\mathcal{G}$ for an arbitrary generator.
Figure 1: Overview of the proposed method.
2.2 D-OPTIMAL SAMPLING STRATEGY
We introduce a graph signal sampling method utilized for our dynamic sensor placement [29]. Here, we define a node-wise sampling operator as follows:
Definition 1 (Node-wise graph signal sampling).
Let $\mathbf{I}_{\mathcal{MV}}\in\{0,1\}^{K\times N}$ be the submatrix of the identity matrix indexed by $\mathcal{M}\subset\mathcal{V}$ $(|\mathcal{M}|=K)$ and $\mathcal{V}$. The sampling operator is defined as
$\mathbf{S}^{\top}\coloneqq\mathbf{I}_{\mathcal{MV}}\mathbf{G},$ (6)
where $\mathbf{G}\in\mathbb{R}^{N\times N}$ is an arbitrary graph filter.
While $\mathbf{G}$ is arbitrary, a typical selection of $\mathbf{G}$ is the identity matrix or a graph lowpass filter as that in classical sampling theory. In a sensor placement perspective, $\mathbf{I}_{\mathcal{MV}}$ specifies the sensor positions on $\mathcal{G}$ and $\mathbf{G}$ can be viewed as a coverage area of sensors. For example, we consider the following polynomial filter
$[\mathbf{G}]_{n,m}=$ $\sum_{i=0}^{N-1}\sum_{p=0}^{P}\alpha_{p}\lambda^{p}u_{i}[n]u_{i}[m]$
$=$ $\left[\mathbf{U}\left(\sum_{p=0}^{P}\alpha_{p}\mathbf{\Lambda}^{p}\right)\mathbf{U}^{\top}\right]_{n,m},$ (7)
where $\{\alpha_{p}\}$ are arbitrary coefficients. In graph signal processing, (7) is referred to as a $P$-hop localized filter[30], whose nonzero response is limited within $P$-hop neighbors of the target node. This implies that the coverage area of all sensors is limited to $P$-hop.
Sensor placement problem is generally formulated as
$\mathcal{M}^{*}=\mathop{\rm arg~{}max}\limits_{\mathcal{M}\subset\mathcal{V}}e(\mathcal{M}),$ (8)
where $e$ is a properly designed cost function. Typically, $e$ is designed based on the optimal experimental designs [31], such as A-, E-, and D-optimality. D-optimality is one of the popular experimental designs and is widely used in graph signal sampling. The D-optimality is more appropriate than other optimality designs if we aim to recover graph signals from non-ideally sampled signals [29].
In the D-optimal design, the following problem is considered as selecting an optimal sampling set for graph signals [29]:
$\mathcal{M}^{*}=\mathop{\arg\max}_{\mathcal{M}\subset\mathcal{V}}\det(\mathbf{Z}_{\mathcal{M}}),$ (9)
where $\mathbf{Z}=\mathbf{G}\mathbf{A}\mathbf{A}^{\top}\mathbf{G}$ and thus $\mathbf{Z}_{\mathcal{M}}=\mathbf{I}_{\mathcal{MV}}\mathbf{G}\mathbf{A}\mathbf{A}^{\top}\mathbf{G}\mathbf{I}^{\top}_{\mathcal{MV}}$. Recall that $\mathbf{S}^{\top}\mathbf{A}\mathbf{A}^{\top}\mathbf{S}$ is invertible when the DS condition is satisfied. Therefore, the maximization of $\det(\mathbf{Z}_{\mathcal{M}})=\det(\mathbf{S}^{\top}\mathbf{A}\mathbf{A}^{\top}\mathbf{S})=|\det(\mathbf{S}^{\top}\mathbf{A})|^{2}$ leads to the best recovery of graph signals.
The direct maximization of (9) is combinatorial and is practically intractable. Therefore, a greedy method is applied to (9). Suppose that $\text{rank}(\mathbf{Z})\geq K$. In [32], (9) can be converted to a greedy selection as follows:
$y^{*}=$ $\mathop{\arg\max}_{y\in\mathcal{M}^{c}}\det(\mathbf{Z}_{\mathcal{M}\cup\{y\}})$
$=$ $\mathop{\arg\max}_{y\in\mathcal{M}^{c}}\mathbf{Z}_{yy}-\mathbf{Z}_{y\mathcal{M}}(\mathbf{Z}_{\mathcal{M}})^{-1}\mathbf{Z}_{\mathcal{M}y}.$ (10)
The selection is performed so that the best node $y^{*}$ is appended to the existing sampling set $\mathcal{M}$ one by one.
3 DYNAMIC SENSOR PLACEMENT ON GRAPHS
In this section, we propose a dynamic sensor placement problem based on sampling theory on graphs. First, we devise online dictionary learning based on sparse coding. Second, we derive the control law of sensors. Fig. 1 shows the overview of the proposed method. The flow of the proposed method is as follows:
  1. 1.
    The dictionary is sequentially learned from previous observations (Section 33.1).
  2. 2.
    Time-varying graph signals are sampled and reconstructed based on graph sampling theory.
  3. 3.
    Sensor positions are dynamically determined based on the dictionary (Section 33.2).
3.1 ONLINE DICTIONARY LEARNING
In sampling and reconstruction introduced in Section 22.1, $\mathbf{A}$ in (2) is assumed to be fixed. Note that (10) is regarded as a static sensor placement problem since the sensor positions are determined only once. In contrast, we consider the time-varying $\mathbf{A}_{t}$ in this paper where the subscript $t$ denotes the time instance (see Table 1). In this setting, the optimal sensor positions can be changed according to $\mathbf{A}_{t}$.
We infer a time-varying generator $\mathbf{A}_{t}$ from previous $D$ observations
$\mathbf{X}_{t-1}=[\mathbf{x}_{t-D},\dots,\mathbf{x}_{t-1}],$ (11)
in every time instance for dynamic sensor placement.
First, we suppose the measurement model as follows:
$\mathbf{Y}_{t}=f(\mathbf{X}_{t-1})+\mathbf{M},$ (12)
where $f$ is a mapping from $\mathbf{X}_{t-1}$ to $\mathbf{X}_{t}$, i.e., $\mathbf{X}_{t}=f(\mathbf{X}_{t-1})$, and $\mathbf{M}$ is additive white Gaussian noise. That is, every observation is generated from signals in the previous time instance. Although the mapping of the time-evolution, $f$, is generally unknown, we suppose that $f$ is Lipschitzian, i.e., there exists some constant $L$ that satisfies the following inequality.
$\|f(\mathbf{X}_{t-1})-f(\mathbf{X}_{t})\|_{F}\leq L\|\mathbf{X}_{t-1}-\mathbf{X}_{t}\|_{F}.$ (13)
By definition of $f$, (13) can be transformed into
$\|\mathbf{X}_{t}-\mathbf{X}_{t+1}\|_{F}\leq L\|\mathbf{X}_{t-1}-\mathbf{X}_{t}\|_{F}.$ (14)
This assumption implies that signals vary smoothly in time [^1] Note that (14) is necessary to stably learn the dictionary, while we do not explicitly estimate $f$ in the following part..
Next, we consider estimating the generator $\mathbf{A}_{t}$ as a dictionary from (the estimated) $\tilde{\mathbf{X}}_{t-1}$ in (11) based on the assumption in (14) where $\tilde{\cdot}$ corresponds to the estimated vector/matrix.
Since the exact value of $M$ in (3) and (4) may not be known in general, herein, we suppose $M\ll D$, and denote a dictionary and expansion coefficient by $\mathbf{A}_{t}\in\mathbb{R}^{N\times D}$ and $\mathbf{d}_{t}\in\mathbb{R}^{D}$, respectively. Therefore, we seek the optimal $\mathbf{A}_{t}$ such that the number in nonzero values of $\mathbf{d}_{t}$ is close to $M$, i.e., $\mathbf{d}_{t}$ is sparse. This setting is similar to well-studied dictionary learning problems [33, 34].
Let $\mathbf{D}_{t}=[\mathbf{d}_{t-D+1},\dots,\mathbf{d}_{t}]\in\mathbb{R}^{D\times D}$ be the collection of expansion coefficients from time $t-D+1$ to $t$. Then, we can express $\mathbf{X}_{t}=\mathbf{A}_{t}\mathbf{D}_{t}$ (see (2)).
Here, we assume that the graph $\mathcal{G}$, initial generator matrix $\mathbf{A}_{0}$, and expansion coefficients matrix $\mathbf{D}_{0}$ for training are given, where $\mathbf{A}_{0}$ and $\mathbf{D}_{0}$ are arbitrary matrices. Following from (14), we can suppose that $\mathbf{X}_{t}$ is sufficiently close to $\mathbf{X}_{t-1}$. As a result, we formulate the following problem:
$\mathop{\rm arg~{}min}\limits_{\mathbf{A}_{t},\mathbf{D}_{t}}\|\tilde{\mathbf{X}}_{t-1}-\mathbf{A}_{t}\mathbf{D}_{t}\|_{F}^{2}+\mu\|\mathbf{D}_{t}\|_{1}+\eta\|\mathbf{A}_{t}-\tilde{\mathbf{A}}_{t-1}\|_{F}^{2},$ (15)
where $\tilde{\mathbf{X}}_{t-1}=\tilde{\mathbf{A}}_{t-1}\tilde{\mathbf{D}}_{t-1}$, and $\mu$ and $\eta$ are the parameters. The first term in (15) is the data fidelity term for $\mathbf{A}_{t}$. The second and third terms are for the regularization controlling the sparsity of $\mathbf{D}_{t}$ and the temporal variation of $\mathbf{A}_{t}$, respectively.
Note that we need to solve (15) with respect to two variables $\mathbf{A}_{t}$ and $\mathbf{D}_{t}$, where they jointly form a nonconvex optimization. In this paper, we divide (15) into two independent problems with respect to $\mathbf{A}_{t}$ and $\mathbf{D}_{t}$, and solve them alternately similarly to the method in [34].
First, we solve (15) with respect to $\mathbf{A}_{t}$ by fixing $\mathbf{D}_{t}$ as $\tilde{\mathbf{D}}_{t-1}$. In this case, we easily obtain the closed-form solution for the dictionary as follows:
$\mathbf{A}_{t}$ $=\mathop{\rm arg~{}min}\limits_{\mathbf{A}_{t}}\|\tilde{\mathbf{X}}_{t-1}-\mathbf{A}_{t}\tilde{\mathbf{D}}_{t-1}\|_{F}^{2}+\eta\|\mathbf{A}_{t}-\tilde{\mathbf{A}}_{t-1}\|_{F}^{2}$
$=(\eta\tilde{\mathbf{A}}_{t-1}+\tilde{\mathbf{X}}_{t-1}\tilde{\mathbf{D}}_{t-1}^{\top})(\eta\mathbf{I}_{D}+\tilde{\mathbf{D}}_{t-1}\tilde{\mathbf{D}}_{t-1}^{\top})^{-1}.$ (16)
Second, we update $\mathbf{D}_{t}$ in (15) by using $\mathbf{A}_{t}$ in (16), i.e.,
$\mathbf{D}_{t}=\mathop{\rm arg~{}min}\limits_{\mathbf{D}_{t}}\|\tilde{\mathbf{X}}_{t-1}-\mathbf{A}_{t}\mathbf{D}_{t}\|_{F}^{2}+\mu\|\mathbf{D}_{t}\|_{1}.$ (17)
To solve (17), we utilize the proximal gradient method[35, 36, 37], whose problem is given by the following form.
$\mathbf{D}_{t}=\mathop{\rm arg~{}min}\limits_{\mathbf{D}_{t}}g(\mathbf{D}_{t})+h(\mathbf{D}_{t}),$ (18)
where $g$ is a differentiable function and $h$ is a proximable function[31]. Note that (18) is set to the entire problem being convex. As a result, the optimal solution is obtained by the following update rule.
$\mathbf{D}_{t,n+1}={\rm prox}_{\gamma h}(\mathbf{D}_{t,n}-\gamma\nabla g(\mathbf{D}_{t,n})),$ (19)
where $\gamma$ is the step size, $\nabla g$ is the gradient of $g$, and ${\rm prox}_{\gamma h}$ is defined as
${\rm prox}_{\gamma h}(\mathbf{D})$ $=\mathop{\rm arg~{}min}\limits_{\mathbf{C}}\frac{1}{2}\|\mathbf{C}-\mathbf{D}\|^{2}_{F}+\gamma\mu\|\mathbf{C}\|_{1}$
$=S_{\gamma\mu}(\mathbf{D}),$ (20)
where $S_{\gamma\mu}$ is the soft thresholding operator defined as follows:
$[S_{\gamma\mu}(\mathbf{D})]_{ij}=\begin{cases}\mathbf{D}_{ij}-\gamma\mu,&\mathbf{D}_{ij}\geq\gamma\mu\\ 0,&|\mathbf{D}_{ij}|<\gamma\mu\\ \mathbf{D}_{ij}+\gamma\mu,&\mathbf{D}_{ij}\leq-\gamma\mu.\end{cases}$ (21)
By applying (20) to (19), we have the following iteration with ISTA [38].
$\mathbf{D}_{t,n+1}$ $=S_{\gamma\mu}(\mathbf{D}_{t,n}-\gamma\nabla g(\mathbf{D}_{t,n}))$
$=S_{\gamma\mu}\{\mathbf{D}_{t,n}-2\gamma\mathbf{A}_{t}^{\top}(\mathbf{A}_{t}\mathbf{D}_{t,n}-\tilde{\mathbf{X}}_{t-1})\},$ (22)
We set the step size as $\gamma\leq 1/{\lambda_{\text{max}}(\mathbf{A}_{t}^{\top}\mathbf{A}_{t})}$ according to the convergence condition of ISTA [38]. We iterates (22) until $\|\mathbf{D}_{t,n+1}-\mathbf{D}_{t,n}\|_{F}^{2}<\epsilon$ where $\epsilon$ is a small constant. We summarize the optimization of (15) in Algorithm 1.
Input: $\tilde{\mathbf{X}}_{t-1},\tilde{\mathbf{A}}_{t-1},\tilde{\mathbf{D}}_{t-1},t$
if $t=1$ then
       $\mathbf{A}_{1}=\mathbf{A}_{0}$
else
       $\mathbf{A}_{t}=(\eta\tilde{\mathbf{A}}_{t-1}+\tilde{\mathbf{X}}_{t-1}\tilde{\mathbf{D}}_{t-1}^{\top})(\eta\mathbf{I}_{D}+\tilde{\mathbf{D}}_{t-1}\mathbf{D}_{t-1}^{\top})^{-1}$
$\mathbf{D}_{t}=\tilde{\mathbf{D}}_{t-1}$
while $\|\mathbf{D}_{t,n}-\mathbf{D}_{t,n-1}\|_{F}^{2}\geq\epsilon$ do
       $\mathbf{D}_{t,n+1}=S_{\gamma\mu}\{\mathbf{D}_{t,n}-2\gamma\mathbf{A}_{t}^{\top}(\mathbf{A}_{t}\mathbf{D}_{t,n}-\tilde{\mathbf{X}}_{t-1})\}$
       $n\leftarrow n+1$
Output: $\tilde{\mathbf{A}}_{t}$
Algorithm 1 Online Dictionary Learning
In the following, we formulate the dynamic sensor placement based on the learned dictionary. The control law of sensors is also derived based on the sampling strategy introduced in (10).
3.2 DYNAMIC SENSOR PLACEMENT
For brevity, we define the following matrix:
$\mathbf{N}_{t}\coloneqq\tilde{\mathbf{A}}_{t}^{\top}\mathbf{G}.$ (23)
Note that we now assume that $\tilde{\mathbf{A}}_{t}$ is given. With (23), (10) can be rewritten by a dynamic form as [^2] For notation simplicity, we omit $\cdot_{t}$ in $\mathbf{N}_{t}$ hereafter.
$\begin{split}y_{t}^{*}=&\mathop{\rm arg~{}max}\limits_{y\in\mathcal{M}^{c}}\mathbf{N}_{:y}^{\top}\mathbf{N}_{:y}-\mathbf{N}_{:y}^{\top}\mathbf{N}_{:\mathcal{M}}(\mathbf{N}_{:\mathcal{M}}^{\top}\mathbf{N}_{:\mathcal{M}})^{-1}\mathbf{N}_{:\mathcal{M}}^{\top}\mathbf{N}_{:y}\\ =&\mathop{\rm arg~{}max}\limits_{y\in\mathcal{M}^{c}}\|\bm{\nu}_{y}\|^{2}-\|\mathbf{P}_{\mathcal{R}(\mathbf{N}_{:\mathcal{M}})}\bm{\nu}_{y}\|^{2},\end{split}$ (24)
where $\bm{\nu}_{y}=\mathbf{N}_{:y}$ and $\mathbf{P}_{\mathcal{R}(\mathbf{Q})}$ is the orthogonal projection onto $\mathcal{R}(\mathbf{Q})$ (both symbols are defined in Table 1). Since $\mathbf{N}$ can temporally vary, the optimal solution in (24) can also change in every time instance.
Note that (24) is based on a greedy selection introduced in Section 22.2. However, in practice, this optimization may not be efficient for large graphs: The control law with (24) implies that we need a large computational burden if we determine the positions of all sensors. This results in a delay for sensor relocations which should be alleviated. In this paper, instead, we consider selecting the sensor positions independently, i.e., (24) is solved for a sensor independently of the other ones. In the following, we rewrite (24) as a distributed optimization.
Here, we denote the position of the $i$th sensor at the $t$th time instance by $q_{i,t}$. We define the $i$th Voronoi region in the graph $\mathcal{G}$ as follows:
Figure 2: Voronoi diagram on a graph. Nodes with circles represent the current sensor positions. Each colored area represents the Voronoi region corresponding to each sensor.
Input: Sensing matrix $\mathbf{G}$, initial positions of sensors $\mathcal{Q}_{1}$, time duration $T$
for $t=1,\ldots,T$ do
       Update the dictionary $\tilde{\mathbf{A}}_{t}$ with Algorithm 1
       Update the data matrix $\tilde{\mathbf{X}}_{t}$
       Calculate the Voronoi regions $\{\mathcal{W}_{i,t}\}_{i=1,\ldots,K}$
       $\mathbf{N}_{t}\leftarrow\tilde{\mathbf{A}}^{\top}_{t}\mathbf{G}$
       par for $i=1,\ldots,K$
             Calculate $\mathbf{P}_{\mathcal{R}(\mathbf{N}_{:\mathcal{M}_{i,t}})}$
             Calculate the $P$ hop neighbors $\{\mathcal{N}^{P}_{i,t}\}_{i=1,\ldots,K}$
             $q^{*}_{i,t+1}=\mathop{\rm arg~{}max}\limits_{y\in\mathcal{W}_{i,t}\cap\mathcal{N}_{i,t}^{P}}\|\bm{\nu}_{y}\|^{2}-\|\mathbf{P}_{\mathcal{R}(\mathbf{N}_{:\mathcal{M}_{i,t}})}\bm{\nu}_{y}\|^{2}$
      $\mathcal{Q}_{t+1}\leftarrow\{q^{*}_{i,t+1}\}_{i=1,\ldots,K}$
Output: Sensor positions $\{\mathcal{Q}_{t}\}_{t=1,\ldots,T+1}$
Algorithm 2 Dynamic Sensor placement
$\mathcal{W}_{i,t}=\{v\in\mathcal{V}\mid d(v,q_{i,t})<d(v,q_{j,t}),\forall q_{j}\in\mathcal{Q}_{t}\},$ (25)
where $d(v,q_{i,t})$ is the shortest distance between $v$ and $q_{i,t}$, and $\mathcal{Q}_{t}=\{q_{j,t}\}_{j=1,\ldots,K}$ denotes the set of all sensor positions at $t$. Note that, in each $\mathcal{W}_{i,t}$, only one sensor is contained: We illustrate its example in Fig. 2.
By utilizing (25), we can rewrite (24) for seeking the best position of a sensor within its Voronoi region $\mathcal{W}_{i,t}$ as follows:
$q^{*}_{i,t+1}=\mathop{\rm arg~{}max}\limits_{y\in\mathcal{W}_{i,t}}\|\bm{\nu}_{y}\|^{2}-\|\mathbf{P}_{\mathcal{R}(\mathbf{N}_{:\mathcal{M}_{i,t}})}\bm{\nu}_{y}\|^{2}.$ (26)
Since (26) depends only on the Voronoi region of a sensor, all sensor selections in (26) are independent of each other. Therefore, we can select sensors by (26) in a distributed fashion.
When we assume mobile sensors, their movable areas could be restricted to their neighbors. To reflect this, we can further impose a constraint on the Voronoi region where the movable nodes are restricted to the $P$-hop neighbors of the target node. The $P$-hop neighbor of the $i$th sensor (node) is defined as
$\mathcal{N}_{i,t}^{P}=\{v\in\mathcal{V}\mid d(v,q_{i,t})\leq P\}.$ (27)
As a result, the proposed control law of sensors in (26) is further modified as follows:
$q^{*}_{i,t+1}=\mathop{\rm arg~{}max}\limits_{y\in\mathcal{W}_{i,t}\cap\mathcal{N}_{i,t}^{P}}\|\bm{\nu}_{y}\|^{2}-\|\mathbf{P}_{\mathcal{R}(\mathbf{N}_{:\mathcal{M}_{i,t}})}\bm{\nu}_{y}\|^{2}.$ (28)
Algorithm 2 summarizes the proposed dynamic sensor placement method [^3] We utilize the MATLAB notation where we denote the parallel processing in Algorithm 2 as par for..
Note that our main contribution is to formulate dynamic sensor placement as graph signal sampling. We may be able to use other methods as an internal algorithm for dictionary learning like [39] and [40].
4 EXPERIMENTS
In this section, we compare reconstruction errors of time-varying graph signals measured by selected positions of sensors. We perform recovery experiments for synthetic and real datasets.
We suppose that the graph $\mathcal{G}$ is given a priori. The nodes represent the candidates of sensor positions, and the edges indicate the paths where sensors can move. Sampling is performed at every time instance. After sampling, the reconstructed signal at $t$, i.e., $\tilde{\mathbf{x}}_{t}$, is stored as the latest signal collection $\tilde{\mathbf{X}}_{t-1}$ which is used for the dictionary/sensor positions update.
Figure 3: MSE comparison of the reconstructed synthetic data for the noisy case.
Figure 4: Reconstructed graph signals for PC signals at time $t=4,9,13$. Circled nodes represent selected sensor positions.
4.1 SYNTHETIC GRAPH SIGNALS
In this subsection, we perform dynamic sensor placement for synthetic data.
4.1.1 SETUP
We compare the signal reconstruction accuracy of the proposed method with those of the following existing methods.
  1. 1)
    Static1: Static sensor placement with the fixed dictionary learned by the SVD only once at $t=0$ [26]. This method does not change the dictionary over time.
  2. 2)
    Static2: Static sensor placement (similar to Static1) with the non-fixed dictionary learned by the SVD at every time instance.
Since the SVD-based dictionary learning is still widely-used in many sensor placement applications[6], Static1 and Static2 are possible baseline methods.
We also compare the proposed method to the extreme case of the proposed method, i.e., $P=\infty$. In this case, sensors can move anywhere in Voronoi region regardless of the current sensor positions.
In the synthetic case, the recovery experiment is performed for signals on a random sensor graph with $N=256$. The number of sensors is set to $K=8$. We consider the following graph signal models:
Bandlimited (BL) model
[…]
$\mathbf{d}_{t}=\begin{bmatrix}\frac{1}{t}\times\sin(10t+\phi_{i})+1\\ \vdots\\ \frac{1}{t}\times\sin(10t+\phi_{M})+1\end{bmatrix},$ (29)
where $\phi_{i}=\frac{\pi}{M}i$. In this model, graph signals temporally oscillate with the fixed bandwidth, and their oscillation amplitudes decay over time. It would mimic a possible physical data observation model.
Piecewise-constant (PC) model
The PC model is defined in (4). The number of clusters is set to $M=3$. We randomly separate nodes into $M$ clusters so that nodes in a cluster are connected. We generate $\mathbf{d}_{t}$ as follows:
$\mathbf{d}_{t}=\begin{bmatrix}3\{\exp({-\frac{t}{25}})\sin(t+\frac{\pi}{3})+1\}\\ 3\{\exp({-\frac{t}{25}})\sin(2t)+1\}\\ 3\{\exp({-\frac{t}{25}})\sin(3t-\frac{\pi}{3})+1\}\end{bmatrix}.$ (30)
In this model, signal values in a cluster change simultaneously (but the temporal frequencies are different for different clusters). It can be a possible model for clustered sensor networks.
For both graph signal models, we generate the data matrix $\mathbf{X}_{t}$ with $T=40$. The number of time slots is set to $D=20$.
While the initial expansion coefficients and dictionary can be arbitrary as mentioned in Section 33.1. In the experiment, however, we used an SVD-based dictionary for a fair comparison to alternative methods. Therefore, $\mathbf{A}_{0}=\bm{\mathcal{U}}$ at $t=0$, where $\bm{\mathcal{U}}$ is obtained by the SVD of the known measurement data, i.e., $\mathbf{X}_{0}=\bm{\mathcal{U}}\mathbf{\Sigma}\bm{\mathcal{V}}^{\top}$. The initial value of $\mathbf{D}_{t}$ is also set to $\mathbf{D}_{0}=\mathbf{11}^{\top}$. We set the temporal sampling period $T_{s}=\frac{\pi}{30}$ to avoid aliasing. The parameters of the dictionary learning are experimentally set to $(\gamma,\eta,\mu)=(10^{-4},3,1)$, and $P=1$ is used for the sensor movable area where we assume a sensor can only be moved to its one-hop neighbor based on reasonable sensor mobility.
We assume that $\mathbf{G}=\mathbf{U}\zeta(\bm{\Lambda})\mathbf{U}^{\top}$ is a graph low-pass filter. We define $\zeta(\bm{\Lambda})=\operatorname{diag}\left(\zeta\left(\lambda_{0}\right),\zeta\left(\lambda_{1}\right),\ldots,\zeta\left(\lambda_{N-1}\right)\right)$ as follows:
$\zeta(\lambda_{i})=\cos\left(\frac{\pi}{2}\cdot\frac{\lambda_{i}}{\lambda_{\text{max}}}\right).$ (31)
Recovery accuracy is measured in noisy cases where white Gaussian noise $w[i]\sim\mathcal{N}(0,0.1)$ is added to the signals. For all methods, the initial position of sensors is determined by a state-of-the-art sampling set selection method on graphs [8]. We calculate the averaged MSEs for 50 independent runs.
Figure 5: Comparison of reconstruction MSEs of sea surface temperature data. The average MSE for each of the methods is plotted as a horizontal line.
Figure 6: Reconstructed errors of sea surface temperature data. Circled nodes represent sensor positions.
4.1.2 RESULTS
Fig. 3 shows the experimental results for the noisy case. We also show examples of the original and reconstructed graph signals of PC models in Fig. 4.
From Fig. 3, the proposed method shows consistently (and significantly) smaller MSEs than the other methods in all time instances, and the MSE of the proposed method gradually decreases over time. This indicates that the proposed method can adapt to the change in measurement conditions, while the static methods fail to do so.
In this experiment, $P=1$ and $\infty$ do not results in a large difference. However, as shown in Fig. 4 at $t=13$, we observe that $P=\infty$ produces a larger error than $P=1$ because several sensor locations are very close. Its possible reason is that the sensor positions were biased on the right side.
4.2 REAL-WORLD SIGNALS
In this subsection, we perform dynamic sensor placement for real data to validate the proposed method for practical observations.
4.2.1 SETUP
We use the global sea surface temperature data [41] during 2016–2021. This dataset is composed of snapshots recorded every month from 2016 to 2021, i.e., its duration is $T=60$. Temperature is sampled at intersections of 1-degree latitude-longitude grids. For simplicity, we clip snapshots of the west coast of the U.S. and randomly sample 100 nodes for the experiments. Then, we construct a $5$-nearest neighbor graph. The numbers of time slots and sensors are set to $D=5$ and $K=10$, respectively. The parameters of the dictionary learning are set to $(\gamma,\eta,\mu)=(10^{-3},1,1)$, and $P=1$ is used for the sensor movable area. We design the filter matrix $\mathbf{G}$ the same as (31) in Section 44.1. We evaluate the averaged MSEs and compare them with the above-mentioned two benchmark methods Static1 and 2.
4.2.2 RESULTS
Fig. 5 shows the MSE comparison. As with the experiment for the synthetic data, the proposed dynamic sensor placement presents lower MSEs than the SVD-based methods for most $t$s. The MSEs of Static1 and Static2 are almost the same.
The reconstruction errors between the original and reconstructed graph signals are visualized in Fig. 6, along with the selected sensor positions. The proposed methods place many sensors in the southern area in March and June, while the positions are more uniformly distributed in September and December. This may reflect the fact that sea surface temperatures in the northern region do not change much from December to April, while those in the southern region change significantly, as shown in Fig. 6. Therefore, in order to capture such changes in sea surface temperatures, the sensors gradually moved to the south after December.
5 CONCLUSION
In this paper, we propose a dynamic sensor placement method based on graph sampling theory. We sequentially learn the dictionary from a time series of observed graph signals by utilizing sparse coding. Using the dictionary, we dynamically determine the sensor placement at every time instance such that the non-observed graph signal values can be best recovered from those of the observed (selected) nodes. In experiments, we demonstrate that the proposed method outperforms existing static sensor placement methods in synthetic and real datasets.